OpenStudy (anonymous):

9^(x^2) 3^7 x = 81 solve for x

OpenStudy (phoenixfire):

\[\large 9^{x^2}+3^{7x}=81\] Is that what you mean?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Recognize that 9, and 81 are all multiples of 3 and you can simply factor that 3 out. Do you understand?

OpenStudy (anonymous):

kind of what would we do

OpenStudy (anonymous):

x^2 + 7 x = 9?

OpenStudy (anonymous):

Close. Shouls be 2x^2 + 7x = 4

OpenStudy (anonymous):

then solve for x?

OpenStudy (anonymous):

wats next

OpenStudy (anonymous):

Yes sir solve for x. its a quadratic equation now, so it should be easy

OpenStudy (anonymous):

how do you get rid of the 2^(x^2) tho?

OpenStudy (anonymous):

Multiply through the equation by 1/2

OpenStudy (anonymous):

So that your equation looks like x^2 + 7x/2 - 2 = 0

OpenStudy (phoenixfire):

@incomplte it's to the power of....\(\large 9^{x^2}\)

OpenStudy (anonymous):

So? @PhoenixFire

OpenStudy (phoenixfire):

\[\large 9^{x^2}+3^{7x}=81\]Dividing through by 3 would give\[\large 3^{x^2}+1^{7x}=27\] I don't really see how that's helped. How do you get the 'x' out of the power to solve it?

OpenStudy (anonymous):

We are combining powers, not dividing by 3

OpenStudy (anonymous):

Remember your basic exponential rules?!

OpenStudy (phoenixfire):

How are you combining powers if the base isn't the same? I'm actually curious as to how this is solved.... My brain doesn't agree :(

OpenStudy (anonymous):

The bases are the same lol

OpenStudy (anonymous):

3^(2x^2)+3^(7x)=3^4

OpenStudy (phoenixfire):

Oh wow.... then \[\large 3^{2x^2+7x}=3^4\]Then take the log to drop the quadratic and the 4 down, the log(3) will cancel on both sides, then you simply solve \[2x^2+7x=4\]

OpenStudy (phoenixfire):

Thanks @incomplte Sorry for the moment of stupidity :P

OpenStudy (anonymous):

No problems